# KEPLA: A Knowledge-Enhanced Deep Learning Framework for Accurate Protein-Ligand Binding Affinity Prediction

### v1

@@ Methods
In parallel with KG embedding, we employ a cross attention network to capture fine-grained, pairwise local interactions between protein and ligand substructures. This module comprises two main steps: (1) constructing an interaction map to measure substructure-level similarity, and (2) applying cross attention to derive a joint protein-ligand representation.
For the first step, we employ separate ESM and GCN encoders to generate hidden representations for protein and ligand: $\mathbf{H}_{p}=\{\mathbf{h}_{p}^{1},\mathbf{h}_{p}^{2},...,\mathbf{h}_{p}^{M}\}$ and $\mathbf{H}_{d}=\{\mathbf{h}_{d}^{1},\mathbf{h}_{d}^{2},...,\mathbf{h}_{d}^{N}\}$, where $M$ and $N$ are the number of encoded subsequences in a protein and the number of atoms in a ligand, respectively. Using these hidden representations, we construct a pairwise interaction matrix $\mathbf{V}\in\mathbb{R}^{M\times N}$:
@@
Each element in $\mathbf{V}$ reflects the interaction strength between a specific protein subsequence and a ligand atom, thus identifying potential binding regions.
To extract a joint representation informed by these local interactions, we treat the interaction matrix as a similarity matrix and compute attention weights over both protein and ligand substructures. Protein attention is calculated by summing the interactions across all ligand atoms and normalizing via a softmax function:
@@
where $\text{Sum}(\cdot)$ denotes column-wise summation in matrix $\mathbf{V}$, and the result accumulates each subsequence’s interactions across ligand atoms. $\sigma(\cdot)$ selects the $\text{Tanh}(\cdot)$ function, and the $\text{Softmax}(\cdot)$ normalizes the attention scores into a probability distribution. The scaling factor $\sqrt{D}$ helps stabilize gradients. Ligand attention is computed analogously by transposing the interaction matrix and aggregating over protein subsequences:
@@
where $\mathbf{f}_{p,d}\in\mathbb{R}^{2D}$ is the joint protein-ligand representation, and $\oplus$ denotes vector concatenation. We feed this joint representation into a multi-layer perceptron decoder to predict the binding affinity:
@@
To jointly learn structural, semantic, and interaction-based knowledge, we combine the affinity prediction objective with the KG embedding loss. The overall training objective is as follows:
$\mathcal{L}=\mathcal{L}_{PLA}+\beta\mathcal{L}_{KE}+\lambda||\boldsymbol{\Theta}||_{2}^{2},$ (17)
@@
We evaluate all models using the following publicly available benchmark datasets for protein-ligand binding affinity prediction. PDBbind is a widely used dataset that provides 3D binding structures of protein-ligand complexes along with experimentally determined binding affinities. In our experiments, we primarily use PDBbind v2016, which is the most frequently adopted version in recent studies. This version consists of three overlapping subsets: The general set, containing all 13,283 protein-ligand complexes; The refined set, comprising 4,057 higher-quality complexes selected from the general set; The core set, including 290 top-quality complexes curated through a rigorous selection process, serving as a benchmark for testing. For convenience, we refer to the 3,767 complexes that belong to the refined set but are not in the core set as the refined set in this paper. To further assess generalization ability, we additionally evaluate on the CSAR-HiQ dataset, an external benchmark consisting of two subsets with 176 and 167 protein-ligand complexes, respectively, sourced independently from PDBbind.
@@
KEPLA is implemented using Python 3.8 and PyTorch 1.7.1 [55], along with functionalities from several auxiliary libraries, including DGL 0.7.1 [56], DGL-LifeSci 0.2.8 [51], Scikit-learn 1.0.2 [57], NumPy 1.20.2 [58], Pandas 1.2.4 [59], and RDKit 2021.03.2 [60]. We set the batch size to 64 and optimize the model using the Adam optimizer with a learning rate of 1e-4. All models are trained for a maximum of 200 epochs, and the model with the lowest validation RMSE is selected for final evaluation on the test set. The protein encoder includes a pre-trained ESM2_t36_3B model, followed by two fully connected layers with dimensions [2560, 512, 128]. The average pooling window size is set to 9. The ligand encoder consists of three GCN layers with hidden dimensions of [128, 128, 128]. We cap the maximum allowed protein sequence length at 1,080 and the maximum number of ligand atoms at 290. In the cross attention module, we use a single attention head for simplicity. The latent embedding size is set to 128, and the fully connected decoder contains 512 hidden units.

### v2

@@ Methods
In parallel with KG embedding, we employ a cross attention module to capture fine-grained, pairwise local interactions between protein and ligand substructures. This module comprises two main steps: (1) constructing an interaction map to measure substructure-level similarity, and (2) applying cross attention to derive a joint protein-ligand representation.
For the first step, ESM and GCN encoders have already generated hidden representations for protein and ligand: $\mathbf{H}_{p}=\{\mathbf{h}_{p}^{1},\mathbf{h}_{p}^{2},...,\mathbf{h}_{p}^{M}\}$ and $\mathbf{H}_{d}=\{\mathbf{h}_{d}^{1},\mathbf{h}_{d}^{2},...,\mathbf{h}_{d}^{N}\}$, where $M$ and $N$ are the number of fragments in a protein and the number of atoms in a ligand, respectively. Using these hidden representations, we construct a pairwise interaction matrix $\mathbf{V}\in\mathbb{R}^{M\times N}$:
@@
Each element in $\mathbf{V}$ reflects the interaction strength between a specific protein fragment and a ligand atom, thus identifying potential binding regions.
To extract a joint representation informed by these local interactions, we treat the interaction matrix as a similarity matrix and compute cross attention weights over both protein and ligand substructures. Protein attention $\boldsymbol{\alpha}_{p}\in\mathbb{R}^{M}$ is calculated by summing the interactions across all ligand atoms and normalizing via a Softmax function:
@@
where $\text{Sum}(\cdot)$ denotes column-wise summation in matrix $\mathbf{V}$, and the result accumulates each fragment’s interactions across ligand atoms. $\sigma(\cdot)$ selects the Tanh function, and the $\text{Softmax}(\cdot)$ normalizes the attention scores into a probability distribution. The scaling factor $\sqrt{D}$ helps stabilize gradients. Ligand attention $\boldsymbol{\alpha}_{d}\in\mathbb{R}^{N}$ is computed analogously by transposing the interaction matrix and aggregating over protein fragments:
@@
where $\mathbf{f}_{p,d}\in\mathbb{R}^{2D}$ is the joint protein-ligand representation, and $\oplus$ denotes vector concatenation. We feed this joint representation into an MLP decoder to predict the binding affinity:
@@
To jointly learn structural and semantic knowledge, we combine the affinity prediction objective with the KG embedding objective. The overall training loss function to minimize is as follows:
$\mathcal{L}=\mathcal{L}_{PLA}+\beta\mathcal{L}_{KGE}+\lambda||\boldsymbol{\Theta}||_{2}^{2},$ (17)
@@
We evaluate KEPLA and baselines using the following publicly available benchmark datasets for PLA prediction. PDBbind is a widely used dataset that provides multi-dimensional binding structures of protein-ligand complexes along with experimentally determined binding affinities. In our experiments, we primarily use PDBbind v2016, which is the most frequently adopted version in recent studies. This version consists of three overlapping subsets: The general set, containing all 13,283 protein-ligand complexes; The refined set, comprising 4,057 higher-quality complexes selected from the general set; The core set, including 290 top-quality complexes curated through a rigorous selection process, serving as a benchmark for testing. For convenience, we refer to the 3,767 complexes that belong to the refined set but are not in the core set as the refined set in this paper. To further assess generalization ability, we additionally evaluate on the CSAR-HiQ dataset, an external benchmark consisting of two subsets with 176 and 167 protein-ligand complexes, respectively, sourced independently from PDBbind.
@@
KEPLA is implemented using Python 3.10 and PyTorch 2.1.2 [57], along with functionalities from several auxiliary libraries, including DGL 2.4.0 [58], DGL-LifeSci 0.3.2 [53], Scikit-learn 1.5.2 [59], NumPy 1.24.3 [60], Pandas 2.2.2 [61], and RDKit 2024.03.5 [62]. The protein encoder includes a pre-trained ESM2_t36_3B model, followed by two fully connected layers with dimensions [2560, 512, 128]. The average pooling window size is set to 9. The ligand encoder consists of three GCN layers with hidden dimensions of [128, 128, 128]. We cap the maximum allowed protein sequence length at 1,080 and the maximum number of ligand atoms at 290. In the cross attention module, we use a single attention head for simplicity. The latent embedding size is set to 128, and the fully connected decoder contains 512 hidden units. We set the batch size to 64 and optimize the model using the Adam optimizer with a learning rate of 1e-4. All models are trained for a maximum of 200 epochs, and the checkpoint with the lowest validation RMSE is selected for final evaluation on the test set.